### v1

# VCR: Visual Caption Restoration

Abstract
We introduce Visual Caption Restoration (VCR), a novel vision-language task that challenges models to accurately restore partially obscured texts using pixel-level hints within images. This task stems from the observation that text embedded in images is intrinsically different from common visual elements and natural language due to the need to align the modalities of vision, text, and text embedded in images. While numerous works have integrated text embedded in images into visual question-answering tasks, approaches to these tasks generally rely on optical character recognition or masked language modeling, thus reducing the task to mainly text-based processing. However, text-based processing becomes ineffective in VCR as accurate text restoration depends on the combined information from provided images, context, and subtle cues from the tiny exposed areas of masked texts. We develop a pipeline to generate synthetic images for the VCR task using image-caption pairs, with adjustable caption visibility to control the task difficulty. With this pipeline, we construct a dataset for VCR called VCR-wiki using images with captions from Wikipedia, comprising English and Chinese entities in both easy and hard configurations. Our results reveal that current vision language models significantly lag behind human performance in the VCR task, and merely fine-tuning the models on our dataset does not lead to notable improvements. We release VCR-wiki and the data construction code to facilitate future research.
1 Introduction
Recent advancements in large language models, such as ChatGPT [50, 49] and Llama [58], have spurred significant interest and progress in the field of vision-language models. With models like GPT-4V [49] and Llava [37, 38, 39] blending textual and visual information, the intersection of computer vision and natural language processing has become a vibrant research frontier. These integrated models aim to leverage the potential of vision and language modalities to understand and interpret multimedia content more effectively.
Amidst this evolving landscape, we introduce Visual Caption Restoration (VCR), a novel vision-language task designed to challenge existing models uniquely. VCR challenges these models to restore obscured texts within images, a task that demands an intricate synthesis of text, vision, and text embedded in the image. The VCR task is grounded in two key insights: (1) text embedded within images, with its characteristics different from common visual elements, represents a distinct modality that requires careful alignment of vision, textual data, and the structure of written texts, and (2) neuroscience findings that suggest that humans are proficient in recognizing partially occluded objects through sophisticated visual and cognitive processes [57, 51, 59, 16, 34]. By leveraging these insights, VCR seeks to explore how well vision-language models can handle texts embedded within images, aligning visual elements and natural language to mimic human-like multi-modal understanding and recognition.
The Visual Question Answering (VQA) task [3, 61, 46, 54] has been a popular benchmark in assessing how well models align and interpret visual and linguistic information. Traditional VQA approaches, however, predominantly focus on direct queries about visible elements in images and do not address the nuanced relationship between textual content embedded within the image and the overall image context. This gap underscores the limited capabilities of current models in processing integrated visual-textual data, particularly when the textual component, which plays a critical role, is partially obscured or altered.
To address these limitations, our VCR task builds on the premise that effective text restoration from images requires an integrated understanding beyond current VQA benchmarks’ capabilities. For instance, in extreme cases, models rely on existing Optical Character Recognition (OCR) system to extract text from documents [54, 7]. The extracted text is then used as context for generating answers, without a true semantic alignment between the text and the visual elements of the document. This approach, while effective in simple scenarios, falls short in more complex settings where text is intricately woven into the visual narrative of the image.
To develop the VCR task, in this work, we introduce a pipeline for generating synthetic images that allows for manipulation of the visibility of the textual components of the image. This not only enhances the challenge posed by the task, but also provides a scalable way to adjust task difficulty. The resulting dataset, VCR-wiki, comprises 2.11M English data and 346K Chinese data sourced from Wikipedia, featuring captions in both languages across ‘easy’ and ‘hard’ difficulty levels. Our evaluations indicate that existing vision-language models significantly underperform compared to human benchmarks, underscoring the need for novel model architectures and training paradigms specifically geared towards this complex intermodal alignment.
By releasing VCR-wiki and the accompanying dataset construction code, we aim to stimulate further research in this area, encouraging the development of models that can more adeptly navigate the nuanced landscape of the restoration of text embedded in images. This effort aligns with the broader goal of advancing vision-language models to achieve deeper, more intuitive understanding of multimedia content, bridging the gap between human and machine perception. Code is available at https://github.com/tianyu-z/VCR.
Contributions
The main contributions of this paper are:
C1
Introduce the Visual Caption Restoration (VCR) task to challenge vision-language models to restore occluded texts in images.
C2
Develop a pipeline for generating synthetic images with embedded text that allows for adjusting visibility of such text, thus providing a rich testing environment for VCR.
C3
Create and release VCR-wiki, a dataset with multilingual captions from Wikipedia images, designed to benchmark vision-language models (VLMs) on text restoration tasks.
C4
Conduct empirical evaluations showing significant gaps between current models and human performance on the VCR task. This highlights the effectiveness of VCR for assessing advancements in VLMs, and underscores the necessity for innovative model architectures and training techniques.
2 VCR Task Description
In this section, we compare the VCR task with other existing tasks and aim to answer the following questions:
Q1
What is the difference between VCR and other visual reconstruction tasks?
Q2
Why should we care about VCR?
For better clarity, we define text embedded in image () as text incorporated within the image. The term visual image (VI) pertains to the portion of the image that excludes the text embedded in the image. The string text (ST) is not part of the image itself, but is associated with it as a distinct textual element. It is usually the question prompt in the form of natural language, e.g., ‘What are the covered texts in the image? Please only guess the covered texts without outputting an explanation.’. Consequently, an element of a Visual Caption Restoration (VCR) task can be expressed as , where is represented as string and both and are presented in image form. This notation does not imply that and can be physically separated into two distinct image components. Instead, this definition is adopted merely to facilitate a clearer explanation of the concepts involved. Please refer to Figure 2 for an illustration of , , and .
A1 Existing tasks that are similar to VCR are the tasks of VQA and OCR. VQA takes as input images and a natural language question, and generates a free-form response. As the ground-truth response is not unique, evaluating VQA poses a major challenge. In contrast to VQA, OCR is a task where the ground truth responses are unique: OCR takes as input complete characters in image form and outputs a string representing the characters in the image, without considering the image context. Models pretrained with OCR are able to retrieve texts embedded in the input image, even if they are incomplete or vague. However, as the vagueness or occlusion of the textual components of the image increases, retrieving the original text without considering the remaining non-textual image context becomes harder, and OCR is no longer a good approach. VCR bridges the gap between OCR and VQA: it reconstructs the unique text found in the image while also considering the visual context of the rest of the image.
Figure 2 is an example VCR task in the hard mode, and Figure 1 shows an example VCR task in the easy mode. Although humans can still fill the blanks easily in the hard mode, it is nearly impossible for models with only OCR capabilities to recover the covered texts without using the context. This is because the pixel-level hints of single characters no longer correspond to a unique solution.
A2 The proposed VCR task is significant in two aspects.
The first aspect of importance stems from neuroscience discoveries about human cognitive abilities to recognize partially occluded objects [16, 34]. Although existing models can recognize objects and texts in images, they often struggle with the complexity of occluded objects due to significant information loss in the images. In contrast, humans excel at filling in missing information using a combination of low-level visual processing and high-level cognitive functions, such as those managed by the prefrontal cortex. This cortical area is known to handle complex cognitive processes like decision-making and memory retention, essential for integrating fragmented visual inputs into coherent objects. We believe that the occlusion restoration task serves as a probe that can effectively distinguish a model’s low-level recognition and high-level cognition involving reasoning. Besides, understanding these neural mechanisms can inspire new algorithms capable of mimicking human-like perception and interpretation in dynamic, real-world conditions where occlusion is common.
The second aspect underscores the unique challenge presented by the VCR task, distinguishing it significantly from existing benchmarks like traditional VQA or occluded object restoration task. By occluding texts instead of common visual objects, VCR targets the models’ text-image alignment capability, which is one of the major challenges for vision-language models. VCR mandates that models align textual and visual information in a manner that replicates human-like understanding involving the utilization of both textual and visual clues. This task necessitates a deep integration of visual (), embedded textual (), and contextual interpretation across modalities, pushing beyond simple text extraction as performed in OCR tasks. In OCR, the focus is primarily on recognizing visible characters, often without a need to understand their contextual relevance within the image narrative. In contrast, VCR introduces complexity by requiring the model to use available partial texts and the visual context collaboratively to reconstruct the obscured content accurately. This not only tests the model’s ability to process - modalities effectively, but also challenges it to maintain internal consistency, akin to human cognitive processes where context and visual clues guide understanding and response. Besides, the difficulty of the task can be adjusted by altering the extent of text occlusion, offering a scalable and flexible framework for systematically enhancing model capabilities in text-visual alignment and semantic comprehension. This rigorous testing ground will help evolve vision-language models to better grasp the nuanced interplay between text and imagery.
3 Dataset Creation
The VCR task requires aligning visual images () with text embedded in images (). Therefore, the dataset creation process relies on a set of highly correlated image-text pairs. We utilize the primary images and their corresponding captions from Wikipedia as the data source111Datasource: https://huggingface.co/datasets/wikimedia/wit_base. to create VCR-wiki, a Wikipedia-based VCR dataset. The creation pipeline for VCR-wiki is shown in Figure 2. Before constructing the dataset, we first filter out the instances with sensitive content, including NSFW and crime-related terms, to mitigate AI risk and biases.
The VCR-wiki dataset is formatted as a VQA task, where each instance includes an image, a question, and a ground-truth answer. The images are synthesized from the text-image pairs by stacking the image () with its corresponding text description () vertically, mimicking the format of a captioned image. This stacked image is referred to as a stacked + image. Each + image is resized to a width of 300 pixels. To prevent excessive image height, we truncate to a maximum of five lines. We filter the dataset to exclude instances with + images exceeding 900 pixels in height to avoid drastic resolution change during data preprocessing.
Within , we use spaCy to randomly select several 5-grams in the caption for masking. To ensure the restoration process is doable by a human without too much professional domain knowledge, the 5-grams do not contain numbers, person names, religious or political groups, facilities, organizations, locations, dates and time labeled by spaCy, and the total masked token does not exceed 50% of the tokens in the caption. We exclude all instances that do not have a single eligible 5-gram. The selected 5-grams are partially obscured by a white rectangle that reveals only the top and bottom parts of the text, with the proportion of coverage varying to adjust task difficulty. Furthermore, to assess the impact of on model performance, we create an ablation for each image, maintaining the resolution of the + image, but retaining only the part in the center of the image.
The VCR task involves a predefined question that prompts the model to produce the text obscured in the image. The ground-truth answer corresponds to the caption displayed in the uncovered portion of the stacked image. Due to the extensive availability of vision-language models and a significant user base in both English and Chinese, we have chosen to develop the dataset in these two languages. For each language, we meticulously select the covering proportion to create two task variants: (1) an easy version, where the task is easy for native speakers but open-source OCR models almost always fail, and (2) a hard version, where the revealed part consists of only one to two pixels for the majority of letters or characters, yet the restoration task remains feasible for native speakers of the language.
We release the dataset under the CC BY-SA 4.0 license. All configurations of VCR-wiki are publicly available at https://huggingface.co/collections/vcr-org/vcr-visual-caption-recognition-6661393b1761e2aff7b967b9.
3.1 Dataset Format and Statistics
| # Train | # Val | # Test | + Image Height | # Obscured Text Spans | |||||||
| Mean | SD | Mean | SD | ||||||||
| English | 2095733 | 5000 | 5000 | 375.52 | 106.01 | 253 | 564 | 1.62 | 0.63 | 1 | 3 |
| Chinese | 336448 | 5000 | 5000 | 360.44 | 102.76 | 239 | 562 | 2.06 | 0.94 | 1 | 4 |
The VCR dataset comprises four configurations: English Easy, English Hard, Chinese Easy, and Chinese Hard. Each configuration can be further divided into training, validation, and test splits. Validation and test splits contain 5,000 entities each. The training set for English configurations and the Chinese configurations contain and instances, respectively, which can be used for model continuous pretraining. We include more detailed statistics of the dataset in Table 1.
4 Experiments
In this section, we report the experiment results of existing state-of-the-art vision-language models on our proposed VCR tasks. The fine-tuning and evaluation of open-source models are conducted on a mix of NVIDIA A100 80G and L40S 48G GPUs in an internal cluster.
4.1 Models
We report evaluation results of the following models:
Closed-source Models.
Open-source Models.
We evaluate open-source models with the best performance on the OpenVLM Leaderboard222We selected the highest-performing open-source models with fewer than 40 billion parameters from the OpenVLM Leaderboard as of May 23, 2024. Details are available at: https://huggingface.co/spaces/opencompass/open_vlm_leaderboard. and state-of-the-art Chinese VLM models. The evaluated models include InternVL-Chat-V1.5[11], MiniCPM-Llama3-V2.5 [25], InternLM-XComposer2-VL-7B [14], CogVLM2-Llama3-19B-Chat [63], Idefics2-8B [32], Yi-VL-34B [1], Yi-VL-6B [1], Qwen-VL-Chat [4], DeepSeek-VL-7B-Chat [42], DeepSeek-VL-1.3B-Chat [42], TextMonkey [40] and DocOwl-1.5 [22]. Out of these models, Idefics2-8B is an English-only model, and CogVLM2-Llama3-19B-Chat has its Chinese variant, CogVLM2-Llama3-19B-Chinese-Chat. Please refer to Table 2 for the model specifications.
Finetuned Models.
To test whether VLMs can learn to conduct VCR via fine-tuning, we select two models from the open-sourced models, CogVLM2-Llama3-19B-Chat and MiniCPM-Llama3-V2.5, and fine-tune them on a subset of VCR’s training set.
More specifically, We fine-tune CogVLM2-Llama3-19B-Chat and MiniCPM-Llama3-V2.5 on the English Hard configuration, and CogVLM2-Llama3-19B-Chinese-Chat and MiniCPM-Llama3-V2.5 on the Chinese Hard configuration. The models are finetuned using LoRA [23] with and . We adopt the schedule-free AdamW optimizer [12] with a learning rate . The effective batch size is 64. Each model is trained on the first 16000 examples of the training set for 1 epoch. All fine-tuning experiments are performed on a single node with 4 NVIDIA L40S 48G GPUs.
4.2 Metrics
We measure the quality of the model’s restoration of each masked -gram (where in our setting, as specified in Section 3). Due to the variability of different models’ outputs, for each masked -gram , where is the vocabulary of the evaluation tokenizer333We utilize spaCy’s en_core_web_sm’s and zh_core_web_sm’s tokenizer for English and Chinese evaluation, respectively., we extract the most similar -gram with the least edit distance in the model’s generation.
We report the two metrics below in our experiment section to measure the restoration quality: Exact Match (), which measures whether the restored -gram totally matches the ground-truth ; and Jaccard Index (), which measures the similarity of and as bag-of-words.
Exact Match (), which measures whether the restored -gram totally matches the ground-truth ;
Jaccard Index (), which is a more relaxed metric that measures the similarity of and as bag-of-words.
where represents the set of tokens in .
4.3 Experimental Results
Please refer to the exact match score and the Jaccard-index of the evaluation in Table 4.3.
Open-Source Models.
We evaluate each open-source model based on the whole 5,000 examples in the test set. Note that Idefics2-8B only supports the English task, hence it has no evaluation score on the Chinese task.
Although achieving state-of-the-art performance on the Open VLM leaderboard, almost all the tested models achieve a low exact match accuracy on the English Easy configuration and fail on the other settings. The best open-source model across the 4 configurations (English Easy, English Hard, Chinese Easy and Chinese Hard) is CogVLM2-Llama3-Chat. This might be attributed to its pretraining process and the special architecture. We also notice that VI has negative impact to most models on the exact match scores () which means the image information is not properly utilized. The best performed open-source model, CogVLM2-Llama3-Chat and its finetuned version has positive except the Chinese Hard configuration. This indicates that information from VI could help improve the model performance on VCR.
For different languages, we noticed a large performance drop when testing the model in Chinese configurations, despite the fact that all the models claim to have basic English-Chinese duolingual capabilities. This is somehow surprising, since Chinese characters, due to their logographic nature, may exhibit a higher degree of recognizability compared to languages that utilize alphabetic scripts in one order [65, 73].
@|c|c|lC1.5cmcccccc|@
Lang
Mode
Model
Model size
Exact match (%)
Jaccard index (%)
EN Easy Closed-source models
Claude 3 Opus - -15 -10.74
GPT-4 Turbo - -3.2 -3.65
GPT-4o - -3.01 -1.32
GPT-4V - 14.17 11.23
Qwen-VL-Max - -8.74 -5.74
Open-source models
CogVLM2 19B 4.96 1.68
CogVLM2-FT 19B 0.64 0.22
DeepSeek-VL 1.3B -8.04 -5.52
DeepSeek-VL 7B -7.93 -4.7
DocOwl-1.5-Omni 8B -0.71 -1.28
TextMonkey 7B -5.54 -5.22
Idefics2 8B -12.02 -19.03
InternLM-XComposer2-VL 7B 0.24 -1.14
InternVL-V1.5 25.5B -60.41 -35.68
MiniCPM-V2.5 8B -8.25 -9.96
MiniCPM-V2.5-FT 8B -3.67 -3.22
Qwen-VL 7B -2.44 -2.34
Yi-VL 34B -0.79 -2.13
Yi-VL 6B -0.9 -2.22
Hard Closed-source models
Claude 3 Opus - -12.2 -12.48
GPT-4 Turbo - -3.5 -2.14
GPT-4o - -9.22 -5.84
GPT-4V - 10.87 14.56
Qwen-VL-Max - -11.07 -9.01
Open-source models
CogVLM2 19B 20.3 20.3
CogVLM2-FT 19B 11.38 6.76
DeepSeek-VL 1.3B -0.23 0.42
DeepSeek-VL 7B -0.75 -1.3
DocOwl-1.5-Omni 8B 0.01 0.03
TextMonkey 7B -0.48 -0.09
Idefics2 8B -0.29 -2.64
InternLM-XComposer2-VL 7B -0.22 -0.72
InternVL-V1.5 25.5B -4.5 -9.67
MiniCPM-V2.5 8B -0.55 -1.43
MiniCPM-V2.5-FT 8B 0.12 0.38
Qwen-VL 7B -0.32 0.77
Yi-VL 34B 0.02 -1.58
Yi-VL 6B 0.02 -1.46
ZH Easy Closed-source models
Qwen-VL-Max - -3.58 -9.42
Open-source models
CogVLM2-Chinese 19B 2.54 3.91
CogVLM2-Chinese-FT 19B 1.84 1.02
DeepSeek-VL 1.3B 0 3.4
DeepSeek-VL 7B 0 -2.76
DocOwl-1.5-Omni 8B 0 -2.23
TextMonkey 7B -0.82 -2.61
InternLM-XComposer2-VL 7B 0.04 0.04
InternVL-V1.5 25.5B -0.55 4.72
MiniCPM-V2.5 8B -0.95 -4.9
MiniCPM-V2.5-FT 8B -0.49 -1.45
Qwen-VL 7B 0.04 1.15
Yi-VL 34B 0 2.64
Yi-VL 6B 0 2.6
Hard Closed-source models
Qwen-VL-Max - -0.49 -6.89
Open-source models
CogVLM2-Chinese 19B -1.32 -2.16
CogVLM2-Chinese-FT 19B -3.51 -3.61
DeepSeek-VL 1.3B 0 3.24
DeepSeek-VL 7B 0 -2.1
DocOwl-1.5-Omni 8B 0 -2.7
TextMonkey 7B 0.05 -0.32
InternLM-XComposer2-VL 7B -0.02 0.46
InternVL-V1.5 25.5B -0.07 2.19
MiniCPM-V2.5 8B 0.01 -0.5
MiniCPM-V2.5-FT 8B 0.42 2.65
Qwen-VL 7B 0 1.06
Yi-VL 34B 0 2.31
Yi-VL 6B 0 2.12
Moreover, we found that models, such as internlm-xcomposer2, are good at OCR and understanding image documents (as demonstrated by DocOwl 1.5 and TextMonkey) still have the potential to be improved in the VCR task. This highlights the unique and indispensable role of VCR in the current suite of benchmarks. Excelling in other document-related benchmarks does not guarantee similar performance in VCR tasks, emphasizing VCR’s distinct challenges and value.
Closed-Source Models.
We evaluate every closed-source model with the first 500 examples in the test set. In English tasks, GPT-4o achieves the best scores among the models which have not been finetuned. Even as GPT-4 series and Claude support Chinese, we found that they cannot completely recognize Chinese words embedded in the image even without any occlusion. Thus, we do not test GPT-4 and Claude model series on Chinese tasks.
In the English configurations, closed-source models outperform all open-source models except CogVLM2, which indicates that model scaling might help improve performance on the VCR task. However, compared with the human evaluation results in Section 4.4, we notice a large performance gap, especially in the English Hard configuration. This shows significant room for improvement in the current state-of-the-art models.
Refer to Table A to compare open-source and closed-source models using the same 500 test cases.
4.4 Human Evaluation
We recruited 7 volunteers to conduct human evaluation on a subset of the samples of our datasets. Two out of the seven evaluators are native English speakers, while five are native Chinese speakers who are also fluent in English171717The participants’ TOEFL scores range from 102/120 to 112/120.. All the volunteers have earned post-graduate degrees majoring in one of biology, statistics, computer science, economy, etc. The evaluations were conducted on a voluntary basis and participants received no rewards.
We gave the volunteer the following instructions: (1) We ask the volunteers to focus on the puzzles. Each example in the hard collection may require 30 seconds to 2 minutes of focused attention; and (2) We ask the volunteers to utilize the context rather than directly brute-force the puzzle.
Every sample is solved by at least 3 volunteers. In English, we release the exact match score in 2 splits: all errors counted (All), and only count errors not related to date and person names (Filtered).
| EN Easy (N = 169) | EN Hard (N = 169) | ZH Easy (N = 188) | ZH Hard (N = 188) | |||||
| Mean (%) | SD (%) | Mean (%) | SD (%) | Mean (%) | SD (%) | Mean (%) | SD (%) | |
| All | 96.65 | 0.34 | 91.12 | 1.18 | 98.58 | 0.31 | 91.84 | 0.81 |
| Filtered | 98.62 | 0.34 | 97.63 | 2.13 | 99.47 | 0.00 | 96.63 | 1.11 |
Refer to Table A to compare all models with human evaluation results using the same test cases.
5 Related Work
Masked modeling.
Masked language modeling (MLM) introduced by BERT [13] and its autoregressive counterparts by GPT [8] have been the foundations of pre-training modern natural language processing (NLP) models. Both are trained to predict a masked-out proportion based on the context. Similarly, inspired by BERT and GPT, iGPT [9] uses the input sequences of pixels to predict the next unknown pixels. Masked Autoencoders (MAE) [19] introduce Masked Image Modeling (MIM) by masking out random patches from input images and reconstructing the missing parts in the images in pixel space. The goal is to encourage the model to learn the representations of the image by understanding the context and structure of the visible patches, which helps reconstruct the masked images accurately.
Visual Question Answering (VQA).
Several datasets have been proposed for visual question answering VQA [3, 71, 17, 46]. FVQA [61] and OK-VQA[43] are datasets about knowledge-based visual question answering and contains questions that necessitate the usage of external knowledge resources. CLEVR [30] is a synthetic VQA dataset that mainly focuses on visual reasoning abilities. Recognizing the need to develop VQA models that can understand text, Text-VQA [54, 6, 47, 64] aims to read and reason about texts embedded within images in the context of image-question answering. Several datasets [54, 6, 47] have been developed for the Text-VQA task, such as the TextVQA dataset [54] and the ST-VQA dataset [6] on natural images, the OCR-VQA dataset [46] on book or movie covers, the InfographicVQA [44] dataset on infographics, and the DocVQA dataset [45] on document images.
Vision Language Model.
Vision-language models are designed for tasks that involve understanding and generating content from images and text [56, 39, 31, 32]. For example, models have been developed to combine Llama3 with advanced vision-language processing capabilities to handle complex multimodal tasks [70, 66, 24, 69, 63, 14]. Qwen-VL [4] enhances visual-linguistic representations for more accurate contextual interpretations, while OpenGVLab-InternVL-Chat [11, 10] merges the InternVL framework with interactive chat capabilities. General-purpose models like GPT-4 series models [50, 49], Claude series models [2] have also been adapted for vision-language tasks, demonstrating strong performance in multimodal tasks. Finally, DocLLM [60] specializes in document understanding by integrating visual and textual data to enhance the interpretation and generation of document-related content. These models collectively represent significant advancements in vision-language integration, contributing unique capabilities and enhancements to the understanding and generation of multimodal information.
Optical Character Recognition (OCR).
OCR [48] and its subproblems [21, 55, 52, 15] have been well-studied in the literature in the constrained setting. However, classical OCR methods often cannot perform well on images captured in the wild in an unconstrained setting. Many new methods have been developed for advancing scene-text recognition on camera-captured images [5, 18, 26, 28, 62, 53, 75, 33]. In addition to the detection and recognition of OCR tasks, visual question answering has emerged as an important downstream task in the OCR literature. With the development of Text-VQA, new methods for improving the reading abilities in VQA utilizing OCR have been proposed. For example, LoRRA [54] extends a VQA model Pythia [29] with an OCR module to better handle Text-VQA tasks. TAP [67] incorporates scene texts that are generated from OCR engines during pretraining to further improve Text-VQA capabilities.
Scene Text Detection (STD).
Identifying and interpreting text in natural images serves as a foundational step towards developing a comprehensive VQA system. This capability allows for the integration of textual information, enhancing overall scene comprehension. Existing approaches mainly consist of two steps: text detection and recognition. Several methods based on Fully Convolutional Neural Networks have been proposed for scene text detection [35, 36, 75, 20, 74]. In terms of text recognition, [27] proposed to directly recognize texts from entire input images using a 90k-class convolutional neural network with each class corresponding to an English word [36]. In addition, several methods [68, 72, 41] have been proposed to tackle the issue of irregular text in scene text detection.
6 Conclusion
In this work, we introduced the Visual Caption Restoration (VCR) task, a novel vision-language challenge aimed at promoting the integration of visual and textual modalities, including text embedded in both natural language tokens and image formats and highly obscured text embedded in the image. We developed a specialized pipeline to create a dataset tailored to this task, utilizing correlated image-text pairs. This task stands out from existing methods by requiring a more profound integration of visual cues and partially obscured text, highlighting its uniqueness and importance in the field.
We conducted extensive evaluations of state-of-the-art vision-language models (VLMs) in both English and Chinese. The results demonstrated significant room for improvement, suggesting that current models have not yet fully harnessed the capabilities necessary for VCR. We selected models representing both the highest and average performance tiers for additional fine-tuning with our dataset. Although fine-tuning exhibited potential for enhancing VCR capabilities, it did not consistently result in significant improvements, indicating the complexity and challenges of adapting models to this task.
By introducing the VCR task and its specialized dataset, we aim to advance research in vision-language interaction. The unique challenges of VCR seek to enhance model development and training, extending the limits of multimodal AI. We invite the community to utilize our dataset and develop innovative strategies to boost VLM model performance.
Acknowledgments and Disclosure of Funding
We acknowledge the contribution of Mila for supporting the evaluation of various models on our datasets through providing GPU access. None of the authors received third-party funding during the last 36 months prior to this submission.
Appendix A Additional evaluation results on first 100 and 500 test cases
@|c|c|lC1.5cmcccccc|@
Lang
Mode
Model
Model size
Exact match (%)
Jaccard index (%)
EN Easy Closed-source models
Claude 3 Opus - -20 -13.06
GPT-4 Turbo - -8.13 -6.07
GPT-4o - -4.78 -1.68
GPT-4V - 7.18 7.15
Qwen-VL-Max - -5.74 -2.82
Open-source models
CogVLM2 19B 1.78 -0.24
CogVLM2-FT 19B -0.59 -0.2
DeepSeek-VL 1.3B -6.51 -4.3
DeepSeek-VL 7B -8.88 -4.01
DocOwl-1.5-Omni 8B -0.59 -0.61
TextMonkey 7B -1.78 -0.76
Idefics2 8B -11.83 -17.76
InternLM-XComposer2-VL 7B 0.59 -0.7
InternVL-V1.5 25.5B -59.76 -33.66
MiniCPM-V2.5 8B -5.92 -5.96
MiniCPM-V2.5-FT 8B -7.69 -6.84
Qwen-VL 7B 0.59 -0.17
Yi-VL 34B 0.59 -1.3
Yi-VL 6B 0.59 -1.81
Hard Closed-source models
Claude 3 Opus - -17 -13.31
GPT-4 Turbo - -4.31 -2.07
GPT-4o - -10.53 -6.21
GPT-4V - 12.44 16.25
Qwen-VL-Max - -14.35 -10.66
Open-source models
CogVLM2 19B 23.67 21.53
CogVLM2-FT 19B 8.28 6.34
DeepSeek-VL 1.3B 0 0.29
DeepSeek-VL 7B -1.18 -1.38
DocOwl-1.5-Omni 8B 0 -0.4
TextMonkey 7B -2.37 -3.31
Idefics2 8B 0.59 -0.53
InternLM-XComposer2-VL 7B -0.59 -1.35
InternVL-V1.5 25.5B -5.33 -10.32
MiniCPM-V2.5 8B -0.59 -0.39
MiniCPM-V2.5-FT 8B -2.96 -1.76
Qwen-VL 7B -1.18 0.63
Yi-VL 34B 0.59 -1.1
Yi-VL 6B 0.59 -0.38
ZH Easy Closed-source models
Qwen-VL-Max - -2.77 -4.97
Open-source models
CogVLM2-Chinese 19B 0.53 1.52
CogVLM2-Chinese-FT 19B -1.06 -2.3
DeepSeek-VL 1.3B 0 3.78
DeepSeek-VL 7B 0 -2.72
DocOwl-1.5-Omni 8B 0 -1.75
TextMonkey 7B 0.53 -3.06
InternLM-XComposer2-VL 7B 0.53 -0.16
InternVL-V1.5 25.5B 1.06 10.59
MiniCPM-V2.5 8B -2.66 -4.81
MiniCPM-V2.5-FT 8B -1.06 -0.66
Qwen-VL 7B 0 0.76
Yi-VL 34B 0 2.69
Yi-VL 6B 0 3.18
Hard Closed-source models
Qwen-VL-Max - -0.79 -4.9
Open-source models
CogVLM2-Chinese 19B 0 -3.05
CogVLM2-Chinese-FT 19B 0.53 -2.95
DeepSeek-VL 1.3B 0 2.34
DeepSeek-VL 7B 0 -2.23
DocOwl-1.5-Omni 8B 0 -2.23
TextMonkey 7B 0 -0.47
InternLM-XComposer2-VL 7B 0 0.18
InternVL-V1.5 25.5B 0 3.03
MiniCPM-V2.5 8B 0 -0.43
MiniCPM-V2.5-FT 8B -1.06 2.03
Qwen-VL 7B 0 1.04
Yi-VL 34B 0 2.15
Yi-VL 6B 0 1.77
@|c|c|lC1.5cmcccccc|@
Lang
Mode
Model
Model size
Exact match (%)
Jaccard index (%)
EN Easy Closed-source models
Claude 3 Opus - -15 -10.74
GPT-4 Turbo - -3.2 -3.65
GPT-4o - -3.01 -1.32
GPT-4V - 14.17 11.23
Qwen-VL-Max - -8.74 -5.74
Open-source models
CogVLM2 19B 3.48 0.79
CogVLM2-FT 19B 0.12 0.06
DeepSeek-VL 1.3B -8.82 -6.62
DeepSeek-VL 7B -7.7 -5.2
DocOwl-1.5-Omni 8B -1.24 -1.44
TextMonkey 7B -6.96 -5.47
Idefics2 8B -14.16 -19.87
InternLM-XComposer2-VL 7B -0.25 -0.65
InternVL-V1.5 25.5B -59.13 -34.82
MiniCPM-V2.5 8B -3.98 -8.32
MiniCPM-V2.5-FT 8B -2.98 -2.46
Qwen-VL 7B -6.71 -4.93
Yi-VL 34B -0.37 -2.02
Yi-VL 6B -0.25 -1.39
Hard Closed-source models
Claude 3 Opus - -12.2 -12.48
GPT-4 Turbo - -3.5 -2.14
GPT-4o - -9.22 -5.84
GPT-4V - 10.87 14.56
Qwen-VL-Max - -11.07 -9.01
Open-source models
CogVLM2 19B 24.97 24.15
CogVLM2-FT 19B 10.68 7.04
DeepSeek-VL 1.3B 0.25 0.01
DeepSeek-VL 7B -0.87 -1.38
DocOwl-1.5-Omni 8B 0 -0.27
TextMonkey 7B -0.87 -1.29
Idefics2 8B 0 -1.75
InternLM-XComposer2-VL 7B 0.12 -0.83
InternVL-V1.5 25.5B -4.6 -9.26
MiniCPM-V2.5 8B 0.12 -0.15
MiniCPM-V2.5-FT 8B -2.86 -1.52
Qwen-VL 7B -0.12 0.85
Yi-VL 34B 0.12 -1.14
Yi-VL 6B 0.12 -1.21
ZH Easy Closed-source models
Qwen-VL-Max - -3.58 -9.42
Open-source models
CogVLM2-Chinese 19B 2.2 3.92
CogVLM2-Chinese-FT 19B 1.3 0.49
DeepSeek-VL 1.3B 0 2.85
DeepSeek-VL 7B 0 -3.02
DocOwl-1.5-Omni 8B 0 -2.63
TextMonkey 7B -1.2 -2.37
InternLM-XComposer2-VL 7B 0.4 -0.18
InternVL-V1.5 25.5B -0.7 5.15
MiniCPM-V2.5 8B -0.3 -4.17
MiniCPM-V2.5-FT 8B 0.2 -1.31
Qwen-VL 7B 0 0.82
Yi-VL 34B 0 2.98
Yi-VL 6B 0 2.62
Hard Closed-source models
Qwen-VL-Max - -0.49 -6.89
Open-source models
CogVLM2-Chinese 19B -1.1 -3.04
CogVLM2-Chinese-FT 19B -3.39 -3.68
DeepSeek-VL 1.3B 0 3.33
DeepSeek-VL 7B 0 -2.08
DocOwl-1.5-Omni 8B 0 -2.73
TextMonkey 7B 0 -0.61
InternLM-XComposer2-VL 7B 0 0.44
InternVL-V1.5 25.5B 0 1.79
MiniCPM-V2.5 8B 0 -0.37
MiniCPM-V2.5-FT 8B -0.2 2.76
Qwen-VL 7B 0 0.94
Yi-VL 34B 0 2.76
Yi-VL 6B 0 1.87
we show the table of evaluation results on first 100 and 500 test cases for better comparison with human evaluation results and closed-source models correspondingly.